# SVBench: A Benchmark with Temporal Multi-Turn Dialogues for Streaming Video Understanding

### v1

@@ 6.4 Performance Analysis of Video Understanding Skills
The analysis of model performance across 9 long-context streaming video understanding skills (see Section 4) in SVBench reveals significant disparities, underscoring the inherent challenges posed by different types of skills. According to Table 3, skills such as Intention Inference (II) and Potentiality Assessment (PA) generally exhibit higher performance across most models, particularly in closed-source LVLMs such as GPT-4V and GPT-4o. This suggests that both open-source and closed-source models are relatively adept at deciphering character intentions and predicting potential future actions. Conversely, more complex skills like Counterfactual Reasoning (CR) and Spatio-Temporal Speculation (STS) manifest comparatively lower accuracy. This trend is echoed in closed-source models, where even the high-performing GPT-4V and GPT-4o models show relative declines in performance, though still maintaining superior scores compared to other models. This drop can be attributed to the cognitive demands of these tasks, which involve abstract reasoning, intricate scenario construction, and temporal-spatial awareness—areas that current models struggle to emulate effectively. RI and ECA also pose significant challenges (see Appendix G). Both involve multi-entity and context-sensitive interactions, leading to performance variability. Notably, StreamingChat outperforms other open-source LVLMs across all 9 skills, and even surpasses closed-source models like GPT in PA, CST, and ECA. Nearly all models scored below 60 on 9 long-context streaming video understanding skills, suggesting that SVBench poses a challenging task.
[…]
tableAblation study on single-instance (Sin.) and multi-turn (Mul.) QA evaluation.
@@ Appendix D Primary Evaluation Metrics
As shown in the Table 4, we employ four primary evaluation metrics to assess the answers generated by various LVLMs mentioned and compare the evaluation results with those evaluated by GPT-4. It can be observed that the results from these four primary evaluation metrics are broadly in agreement with the evaluation made by GPT-4, which suggests that our SVBench is capable of accurately and effectively distinguishing distinct levels of capabilities possessed by various LVLMs. Just like the use of sentence similarity Di & Xie (2024) in QAEgo4D as an important evaluation metric for computing distances in semantic space, we will subsequently use semantic similarity as an evaluation criterion to assess the consistency between the model’s output answers and the ground truth.
@@ F.1 Comparison of LLM-based Evaluations with Human Evaluation
Figure 7: Score comparison on multi-turn QA evaluation (Mul.).
@@ F.3 Additional Evaluation Results
We include a comparison with human performance, as shown in 5, which can provide valuable insights into the gap between current models and human capabilities in long-context streaming video understanding. The results show that human performance significantly outperforms various open-source and closed-source models across all metrics in both evaluation settings (Dialogue Evaluation and Streaming Evaluation). Additionally, humans excel in Temporal Understanding (TU) but perform relatively weaker in Informational Completeness (IC). We also included LVLMs such as Flash-VStream-7B Zhang et al. (2024a), Qwen2-VL-7B Wang et al. (2024a), and LLaVA-NeXT-Video-7B-DPO Zhang et al. (2024c). We will also add the evaluation results for the models (Oryx Liu et al. (2024b), Long-LLaVA Wang et al. (2024c), LongVILA Xue et al. (2024)) to Table 5 as soon as possible.
@@ F.4 Impact of Model Size on Performance
Appendix G Prompts

### v2

@@ 6.4 Performance Analysis of Video Understanding Skills
The analysis of model performance across 9 long-context streaming video understanding skills (see Section 4) in SVBench reveals significant disparities, underscoring the inherent challenges posed by different types of skills. According to Table 3, skills such as Intention Inference (II) and Potentiality Assessment (PA) generally exhibit higher performance across most models, particularly in closed-source LVLMs such as GPT-4V and GPT-4o. This suggests that both open-source and closed-source models are relatively adept at deciphering character intentions and predicting potential future actions. Conversely, more complex skills like Counterfactual Reasoning (CR) and Spatio-Temporal Speculation (STS) manifest comparatively lower accuracy. This trend is echoed in closed-source models, where even the high-performing GPT-4V and GPT-4o models show relative declines in performance, though still maintaining superior scores compared to other models. This drop can be attributed to the cognitive demands of these tasks, which involve abstract reasoning, intricate scenario construction, and temporal-spatial awareness—areas that current models struggle to emulate effectively. RI and ECA also pose significant challenges (see Appendix H). Both involve multi-entity and context-sensitive interactions, leading to performance variability. Notably, StreamingChat outperforms other open-source LVLMs across all 9 skills, and even surpasses closed-source models like GPT in PA, CST, and ECA. Nearly all models scored below 60 on 9 long-context streaming video understanding skills, suggesting that SVBench poses a challenging task.
Table 4: Ablation study on single-instance (Sin.) and multi-turn (Mul.) QA evaluation.
@@ Appendix D Primary Evaluation Metrics
As shown in the Table 5, we employ four primary evaluation metrics to assess the answers generated by various LVLMs mentioned and compare the evaluation results with those evaluated by GPT-4. It can be observed that the results from these four primary evaluation metrics are broadly in agreement with the evaluation made by GPT-4, which suggests that our SVBench is capable of accurately and effectively distinguishing distinct levels of capabilities possessed by various LVLMs. Just like the use of sentence similarity di2024grounded in QAEgo4D as an important evaluation metric for computing distances in semantic space, we will subsequently use semantic similarity as an evaluation criterion to assess the consistency between the model’s output answers and the ground truth.
@@ F.1 Comparison of LLM-based Evaluations with Human Evaluation
Figure 7: Score comparison on multi-turn QA evaluation (Mul.).
@@ F.3 Additional Evaluation Results
We include a comparison with human performance, as shown in 6, which can provide valuable insights into the gap between current models and human capabilities in long-context streaming video understanding. The results show that human performance significantly outperforms various open-source and closed-source models across all metrics in both evaluation settings (Dialogue Evaluation and Streaming Evaluation). Additionally, humans excel in Temporal Understanding (TU) but perform relatively weaker in Informational Completeness (IC). We also included LVLMs such as Flash-VStream-7B zhang2024flash, Qwen2-VL-7B Qwen2VL, and LLaVA-NeXT-Video-7B-DPO zhang2024llavanext-video. We will also add the evaluation results for the models (Oryx liu2024oryx, Long-LLaVA wang2024longllavascalingmultimodalllms, LongVILA longvila) to Table 6 as soon as possible.
@@ F.4 Impact of Model Size on Performance
Appendix G Real-Time Leaderboard
Our benchmark is an ongoing project, and we plan to continuously update it with various mainstream Video-LLMs, maintaining a real-time updated leaderboard. Given the resource-intensive nature of evaluating a large number of newly emerging models using GPT-4o, we have switched to using the latest and best open-source model, Qwen2.5-72B-Instruct, for our assessments and scoring. The real-time updated leaderboard can be found at https://github.com/yzy-bupt/SVBench.
Appendix H Prompts